When two cars A and B are heading towards each other they approach 20m every second .when the two are moving away from each other they approach 60m apart each second . Find the velocity of each car.
1
Expert's answer
2018-12-24T09:58:11-0500

We have:

v_1-v_2=20 m/s.

v_1+v_2=60 m/s.

Thus,

2v_1=20+60=80 m/s

v_1=80/2=40 m/s

And

v_2=60-40=20 m/s.